OpenStudy (anonymous):

Write an appropriate inverse variation equation if y = 5 when x = 3

OpenStudy (anonymous):

inverse variation equation y proportional to 1/x y=k/x k an arbitrary constant xy=k when x=3 y=5 put these xy=15